# Effective Medium Theory for Time-modulated Subwavelength Resonators

### v1

@@ 4.1 Homogenised Equation
for all $i=1,\dots,N$ and $n=-K,\dots,K$. Now we define the vectors $\mathbf{x},\mathbf{b}\in\mathbb{C}^{N(2K+1)}$ with entries
$\mathbf{x}=\begin{bmatrix}v^{\mathrm{in}}_{1,-K}(z_{1})\\ \vdots\\ v^{\mathrm{in}}_{1,K}(z_{1})\\ \vdots\\ v^{\mathrm{in}}_{N,-K}(z_{N})\\ \vdots\\ v^{\mathrm{in}}_{N,K}(z_{N})\end{bmatrix},\quad\mathbf{b}=\begin{bmatrix}v^{\mathrm{in}}_{-K}(z_{1})\\ \vdots\\ v^{\mathrm{in}}_{K}(z_{1})\\ \vdots\\ v^{\mathrm{in}}_{-K}(z_{N})\\ \vdots\\ v^{\mathrm{in}}_{K}(z_{N})\end{bmatrix}$ (4.28)
@@
$T^{ij}_{nm}:=$$\begin{cases}\tilde{g}_{mn}^{j}G^{k^{(m)}}(z_{i}-z_{j}),&i\neq j,\\ 0,&i=j,\end{cases}\qquad T:=\begin{bmatrix}\begin{bmatrix}T^{ij}_{nm}\end{bmatrix}_{n,m=-K}^{K}\end{bmatrix}_{i,j=1}^{N}.$ (4.29)
This notation allow us to rewrite (4.27) as
$\mathbf{x}-T\mathbf{x}=\mathbf{b},$ (4.30)
which is of the same form as in the static case (3.17), but where the linear system has a block structure. As in the static case, we must pose some assumptions in order to derive the homogenised equation.
@@
Assume that $\kappa_{i}(t)\equiv\kappa(t)$ is identical across all resonators. Therefore, $\lambda^{i}_{j}=c^{i}_{j}\ell\equiv c_{j}\ell$ and as a consequence, $g^{i}_{mn}\equiv g_{mn}$.
We also impose the assumptions of the previous section: 3.4 (constant volume fraction), 3.5 (limiting density $\tilde{V}(x)$), and 3.6 (continuum limit approximation). This allow us to write
@@
The definition of $\beta_{nm}$ was obtained in the same way as in the static case (3.18). Recall that $g^{i}_{nm}$ defined by (4.12) does not depend on $\ell$ and therefore, $\beta^{i}_{nm}$ does not depend on $\ell$ nor $N$. Hence,
@@
For any $f_{m}\in C^{0,\alpha}(\mathcal{U})$ with $0<\alpha\leq 1$,
@@
for all $n=-K,\dots,K$.
Proof.
First, define $V_{nm}(x):=\beta_{nm}\Lambda\tilde{V}(x)$, for $\tilde{V}$ defined through (3.21). We use Assumption 3.6 to obtain the desired result. ∎
@@
By substituting the above defined integral operator $\mathcal{T}^{nm}$ into Lemma 4.5, applying it to (4.32), and neglecting the remainder term, we obtain a system of coupled Lippmann-Schwinger equations
$\psi_{n}+\sum_{m=-K}^{K}\mathcal{T}^{nm}\psi_{m}=\psi^{\mathrm{in}},\quad\forall\,n=-K,\dots,K.$ (4.35)
Then, applying the operator $\left(\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\left(k^{(n)}\right)^{2}\right)$ leads to the homogenised equations
$\left(\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\left(k^{(n)}\right)^{2}\right)\psi_{n}(x)-V_{nn}(x)\psi_{n}(x)\\ -\sum\limits_{m=-K}^{K}\left(\left(k^{(m)}\right)^{2}-\left(k^{(n)}\right)^{2}\right)\int_{\mathcal{U}}G^{k^{(n)}}(x-y)V_{nm}(y)\psi_{m}(y)\,\mathrm{d}y=0.$ (4.36)
@@
Note that ${\psi}_{n}(z_{i})$ is the continuum limit of ${v}^{\mathrm{in}}_{i,n}$ as $N\to\infty$, which is proved in detail for the higher-dimensional static case in [5]. Looking at (4.23) and (4.27), the total field is given by $v^{\mathrm{in}}_{i,n}+v^{\mathrm{sc}}_{i,n}=v_{n}$ for all $i=1,\dots,N$. However, as $N\to\infty$, the scattered field of a single resonator $v_{i,n}^{\mathrm{sc}}$ becomes negligible. Thus, ${\psi}_{n}$ is the continuum limit of the total wave field $v_{n}$.
@@
$\mathrm{diag}\left(\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\left(k^{(n)}\right)^{2}\right)_{n=-K}^{K}\boldsymbol{\psi}(x)-\mathrm{diag}\left(V_{nn}(x)\right)_{n=-K}^{K}\boldsymbol{\psi}(x)-\int_{\mathcal{U}}\mathcal{A}(x,y)\boldsymbol{\psi}(y)\,\mathrm{d}y=\mathbf{0},$ (4.37)
for the vector $\boldsymbol{\psi}:=\left(\psi_{n}\right)_{n=-K}^{K}$ and the matrix $\mathcal{A}(x,y):=\left(\mathcal{A}_{nm}(x,y)\right)_{m,n=-K}^{K}$ with entries
$\mathcal{A}_{nm}(x,y):=\begin{cases}0,&n=m,\\ \left(\left(k^{(m)}\right)^{2}-\left(k^{(n)}\right)^{2}\right)G^{k^{(n)}}(x-y)V_{nm}(y),&n\neq m.\end{cases}$ (4.38)
@@
The system of coupled equations (4.35) can be rewritten into a single vector Lippmann-Schwinger equation given by
$\boldsymbol{\psi}(x)=\mathbf{v}^{\mathrm{in}}(x)+\int_{\mathcal{U}}\mathcal{V}(x,y)\boldsymbol{\psi}(y)\,\mathrm{d}y,$ (4.39)
@@
$\mathcal{V}(x,y):=\left(\mathcal{V}_{nm}(x,y)\right)_{n,m=-K}^{K},\quad\mathcal{V}_{nm}(x,y):=G^{k^{(n)}}(x-y)V_{nm}(y).$ (4.40)
4.7 and 4.8 are the main results of the homogenisation. They provide two equivalent formulations of the effective field $\boldsymbol{\psi}$ emerging from a homogenised time-dependent material. Comparing the homogenised equation (4.37) to the static case in (3.24), we observe that the time-modulated case is described by an integro-differential homogenised equation, rather than a differential equation.
@@
As pointed out in [5], we note that it is not possible to obtain an effective medium theory for frequencies that are too close to the resonant frequency of a high-contrast resonator. In this case, the scattering matrix $g$ is of order one as $\ell\to 0$. Each additional resonator therefore yields a strong contribution to the total field. The limit as $N\to\infty$ does not exist and homogenisation is not possible.
To justify the asymptotic regime used in Section 4, we use the formulas for the resonant frequencies of a single resonator proved in [2]:
$\omega_{0}=0,\quad\omega_{1}\approx-\frac{2\mathrm{i}(v_{\mathrm{r}})^{2}\delta}{\ell v_{0}T}\int_{0}^{T}{\kappa(t)}\,\mathrm{d}t,$ (4.41)
as $\delta\to 0$. If $\delta=\gamma\ell^{p}$, for some $p\in\mathbb{N}_{>1}$, having an incident wave at resonance would mean that $\omega=\mu\ell^{p-1}$. However, under this assumption, the term $\beta_{mn}\Lambda v^{\mathrm{in}}_{i,n}(z_{j})G^{k^{(n)}}(z_{i}-z_{j})$ inside (4.32) is of order one as $\ell\to 0$. As $N\to\infty$, this sum will tend to infinity. Therefore, to obtain an effective medium theory, we must work in a regime for which $g=O(\ell)$. The only possible regime is $\omega=O(\delta)$ and $\delta=O(\ell^{q})$ for some $q>0$. For simplicity, we choose $q=1$.
@@ 5 Numerical Results
We now present a modified Nyström numerical scheme (similar to [8]) to solve the system of Lippmann-Schwinger equations (4.39). Let $\left(y_{i}\right)_{i=1}^{M}$ be a set of $M$ uniformly distributed points in $\mathcal{U}$. Discretising the Lippmann-Schwinger equation yields
$\boldsymbol{\psi}(y_{i})=\mathbf{v}^{\mathrm{in}}(y_{i})+\frac{|\mathcal{U}|}{M}\sum\limits_{j=1}^{M}\mathcal{V}(y_{i},y_{j})\boldsymbol{\psi}(y_{j}),$ (5.1)
@@
$\boldsymbol{\psi}^{M}:=\begin{bmatrix}\boldsymbol{\psi}(y_{1})\\ \vdots\\ \boldsymbol{\psi}(y_{M})\end{bmatrix},\quad\mathbf{v}^{\mathrm{in},M}:=\begin{bmatrix}\mathbf{v}^{\mathrm{in}}(y_{1})\\ \vdots\\ \mathbf{v}^{\mathrm{in}}(y_{M})\end{bmatrix},\quad\mathbb{V}:=\left(\mathcal{V}(y_{i},y_{j})\right)_{i,j=1}^{M}.$ (5.2)
@@
$\left(I_{M(2K+1)}-\frac{|\mathcal{U}|}{M}\mathbb{V}\right)\boldsymbol{\psi}^{M}=\mathbf{v}^{\mathrm{in},M}.$ (5.3)
@@
Figure 5.1: The effective field $\psi$ (-) and the total field $u$ (*) plotted next to each other over $\mathcal{U}:=[0,4000]$ with $N=100$ resonators each of length $\ell=0.01$, thus $\Lambda=1$. We set $\gamma=0.05,\,\mu=0.11,\,\xi=0.2,\,K=4,\,\varepsilon_{\kappa}=0.4$. We evaluate the solution at ten time-steps in $[0,T]$, for $T=2\pi/\Omega=3141.59$. For the definition of (5.3) we choose a mesh of size $M=200$.
Figure 5.2: The $\ell^{2}$-norm of the error between the effective field $\psi^{M}$ and the total field $u$ at time $t=0$. For the numerical solution we set $M=N$ and $y_{i}=z_{i}$, for convenience. The resonators are evenly distributed inside $\mathcal{U}=[0,4000]$ and each of length $\ell=1/N$. We set $\gamma=0.05,\,\mu=0.11,\,\xi=0.2,\,K=4,\,\varepsilon_{\kappa}=0.4$.
The bold line in Figure 5.1 shows the effective field $\psi^{M}$ as a function of $x$ evaluated at different time-steps inside the interval $[0,T]$. For comparison, we plot the evaluation of the total wave field obtained using the scheme obtained in [2], marked by stars inside the plot. As analytically proved, the effective field and the total field agree for $\ell\ll 1$ and large $N$. The numerical results in Figure 5.1 show that the effective field is still quasi-periodic with quasi-periodicity $\omega$.
Figure 5.2 shows the $\ell^{2}$-norm of the error between the effective field $\psi$ and the true solution $u$ at each resonator $z_{i}$ as a function of $N$. It becomes apparent from our numerical result that the error behaves algebraically in $N$. This result is not surprising, as the effective medium theory only holds in the limit $\ell\to 0$ and $N\to\infty$.
Figure 5.3: The runtime of the computation of $\psi$ solving (5.3) and $u$ solving (2.6). These results correspond to the same setting as in Figure 5.2.
The results in Figure 5.3 show how the computation run time of $u$ is significantly larger than the computation of the effective field $\psi$. This further underlines the relevance of the herein derived effective medium theory.

### v3

@@ 4.1 Homogenised Model
for all $i=1,\dots,N$ and $n=-K,\dots,K$. Now we define the vectors $\boldsymbol{x},\boldsymbol{b}\in\mathbb{C}^{N(2K+1)}$ with entries
$\boldsymbol{x}=\begin{bmatrix}v^{\mathrm{in}}_{1,-K}(z_{1})\\ \vdots\\ v^{\mathrm{in}}_{1,K}(z_{1})\\ \vdots\\ v^{\mathrm{in}}_{N,-K}(z_{N})\\ \vdots\\ v^{\mathrm{in}}_{N,K}(z_{N})\end{bmatrix},\quad\boldsymbol{b}=\begin{bmatrix}v^{\mathrm{in}}_{-K}(z_{1})\\ \vdots\\ v^{\mathrm{in}}_{K}(z_{1})\\ \vdots\\ v^{\mathrm{in}}_{-K}(z_{N})\\ \vdots\\ v^{\mathrm{in}}_{K}(z_{N})\end{bmatrix}$ (4.30)
@@
$T^{ij}_{nm}:=$$\begin{cases}\tilde{g}_{nm}^{j}G^{k^{(n)}}(z_{i}-z_{j}),&i\neq j,\\ 0,&i=j,\end{cases}\qquad T:=\begin{bmatrix}\begin{bmatrix}T^{ij}_{nm}\end{bmatrix}_{n,m=-K}^{K}\end{bmatrix}_{i,j=1}^{N}.$ (4.31)
This notation allows us to rewrite (4.29) as
$\boldsymbol{x}-T\boldsymbol{x}=\boldsymbol{b},$ (4.32)
which is of the same form as in the static case (3.17), but here the linear system has a block structure. As in the static case, we must pose some assumptions in order to derive the homogenised model.
@@
Assume that $\kappa_{i}(t)\equiv\kappa(t)$ is identical across all resonators. Therefore, $\lambda^{i}_{j}=c^{i}_{j}\ell\equiv c_{j}\ell$ and as a consequence, $g^{i}_{nm}\equiv g_{nm}$.
We also impose the assumptions of the previous section: 3.4 (constant volume fraction), 3.5 (limiting density $\tilde{V}(x)$), and 3.6 (continuum limit approximation). This allows us to write
@@
The definition of $\beta_{nm}$ was obtained in the same way as in the static case (3.18). Recall that $g^{i}_{nm}$ defined by (4.14) does not depend on $\ell$ and therefore $\beta^{i}_{nm}$ does not depend on $\ell$ nor $N$. Hence,
@@
First, we let
$V_{nm}(x):=\begin{cases}\beta_{nm}\Lambda\tilde{V}(x),&x\in\mathcal{U},\\ 0,&x\in\mathbb{R}\setminus\mathcal{U},\end{cases}$ (4.35)
for $\tilde{V}$ defined through (3.21). In view of the definitions (4.33) and (4.35) of $\beta_{nm}$ and $V_{nm}$, and the fact that $K$ is fixed, we can use Assumption 3.6 to obtain the following result.
@@
For any $f_{m}\in C^{0,\alpha}(\mathcal{U})$, $m=-K,\ldots,K,$ with $0<\alpha\leq 1$,
@@
for all $n=-K,\dots,K$. Here, $C_{nm}$ is independent of $N$ and $\{f_{m}\}_{m=-K}^{K}.$
@@
By substituting the above defined integral operator $\mathcal{T}^{nm}$ into Lemma 4.5, applying it to (4.34), and neglecting the remainder term, we obtain the following system of coupled Lippmann-Schwinger equations:
$\psi_{n}-\sum_{m=-K}^{K}\mathcal{T}^{nm}\psi_{m}=\psi^{\mathrm{in}},\quad\forall\,n=-K,\dots,K.$ (4.38)
Then, applying the operator $\left(\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\left(k^{(n)}\right)^{2}\right)$ leads to the coupled system of differential equations
$\left(\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+\left(k^{(n)}\right)^{2}\right)\psi_{n}(x)-\sum\limits_{m=-K}^{K}V_{nm}(x)\psi_{m}(x)=0,\quad x\in\mathbb{R},$ (4.39)
which describes the effective properties of our system of time-modulated subwavelength resonators.
@@
We first emphasise that the term $V_{nm}(x)$, which incorporates the homogenised effect of the time-modulated resonators, vanishes outside of $\mathcal{U}$. Moreover, in (4.39), both $\psi_{n}$ and its derivative are continuous across the boundary of $\mathcal{U}$. Note also that ${\psi}_{n}(z_{i})$ is the continuum limit of ${v}^{\mathrm{in}}_{i,n}$ as $N\to\infty$, which is proved in detail for the higher-dimensional static case in [7]. Looking at (4.25) and (4.29), the total field is given by $v^{\mathrm{in}}_{i,n}+v^{\mathrm{sc}}_{i,n}=v_{n}$ for all $i=1,\dots,N$. However, as $N\to\infty$, the scattered field of a single resonator $v_{i,n}^{\mathrm{sc}}$ becomes negligible. Thus, ${\psi}_{n}$ is the continuum limit of the total wave field $v_{n}$.
@@
$\left(\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}+K^{2}\right)\boldsymbol{\psi}(x)-\mathcal{V}(x)\boldsymbol{\psi}(x)=\boldsymbol{0},$ (4.40)
for $x\in\mathbb{R}$, where the vector $\boldsymbol{\psi}:=\left(\psi_{n}\right)_{n=-K}^{K}$ and the matrices $K$ and $\mathcal{V}$ are given by $K:=\mathrm{diag}\left(k^{(n)}\right)_{n=-K}^{K}$ and $\mathcal{V}(x):=\left(V_{nm}(x)\right)_{m,n=-K}^{K}$.
@@
Assume the material parameters to be scaled as in Lemma 4.2. The system of coupled equations (4.38) can then be rewritten into a single vector Lippmann-Schwinger equation given by
$\boldsymbol{\psi}(x)=\boldsymbol{v}^{\mathrm{in}}(x)+\int_{\mathcal{U}}\mathcal{G}(x-y)\mathcal{V}(y)\boldsymbol{\psi}(y)\,\mathrm{d}y,\quad x\in\mathbb{R},$ (4.41)
@@
$\mathcal{V}(x):=\left({V}_{nm}(x)\right)_{n,m=-K}^{K},\quad\mathcal{G}(x):=\mathrm{diag}\left(G^{k^{(n)}}(x)\right)_{n=-K}^{K}.$ (4.42)
4.7 and 4.8 are the main results of our homogenisation theory for systems of time-modulated subwavelength resonators. They provide two equivalent formulations of the effective field $\boldsymbol{\psi}$. Note that the homogenisation model described by (4.40) holds under Assumptions (3.4), (3.5), (3.6) and (4.4), and is valid in the regime where the operating frequency is of order $\omega=O(\delta)$.
Comparing the homogenised equation (4.40) to the one in the static case in (3.24), we observe that the effective model in the time-modulated case is described by a system of coupled differential equations, rather than by a single differential equation as in the static case. This is due to the fact that time-modulations give rise to a family of coupled harmonics at frequencies $(\omega+n\Omega)/v_{\mathrm{r}}$. Note also that in both effective models, the scattering coefficients $g$ and $g_{nm}$ depend on the operating frequency $\omega$. Moreover, as in Remark (3.7), positivity of the matrix induces a dissipation effect on components of the effective field $\boldsymbol{\psi}$.
@@
To justify the asymptotic regime used in this section, we recall the formulas for the two subwavelength resonant frequencies $\omega_{0},\omega_{1}$ of a single resonator in one dimension proved in [3]:
$\omega_{0}=0,\quad\omega_{1}\approx-\frac{2\mathrm{i}(v_{\mathrm{r}})^{2}\delta}{\ell v_{0}T}\int_{0}^{T}{\kappa(t)}\,\mathrm{d}t+O\left(\frac{\delta^{3/2}}{\ell}\right),$ (4.43)
as $\delta\to 0$. In particular, we note that $\mathrm{Re}\left(\omega_{1}\right)=O(\delta^{3/2}/\ell)$. Choosing $\delta=O(\ell)$ yields $\mathrm{Re}\left(\omega_{1}\right)=O(\delta^{1/2})$. Now, if we consider an operating frequency $\omega$ of order $\mathrm{Re}\left(\omega_{1}\right)$, then the summand of (4.34) is of order one as $\delta\to 0$. As $N\to\infty$, this sum cannot converge and hence it is not possible to obtain an effective medium theory. Therefore, to obtain an effective medium theory, we must work in a regime for which the scattering coefficient $g$ of a single resonator goes to zero as $\delta\to 0$. To achieve this, one simple choice is to take $\omega=O(\delta).$
Remark 4.10.
As in Remark 3.7, it is also worth emphasising that the homogenised model (4.40) may not have a unique solution. Nevertheless, by exactly the same arguments as those in [7], we can prove that if the homogenised model is well-posed, then we have $L^{2}$-convergence of the solution of the scattering problem with $N$ resonators to the solution of (4.40) as $N$ goes to infinity.
Remark 4.11.
Note that in higher dimensions the characterisation (4.25) still holds true. However, the scattered wave field is expressed in terms of the single layer potential $\mathcal{S}^{k}_{D}[\psi]$ through [7]
$v_{i,n}^{\mathrm{sc}}(x):=\mathcal{S}^{k^{(n)}}_{D_{i}}[\psi_{i,n}](x),$ (4.44)
for some surface densities $\psi_{i,n}\in L^{2}(\partial D)$, for all $i=1,\dots,N$ and $n\in\mathbb{Z}$. Then we can proceed in an equivalent manner as in the one-dimensional case to obtain a linear system of equations of the form (4.32). With a set of assumptions similar to those stated here, but with a higher-dimensional notation as in [7, Section 2], we would arrive at a result similar to the one stated in Corollary 4.7.
@@ 5 Numerical Results
We now present a standard Nyström numerical scheme to solve the system of Lippmann-Schwinger equations (4.41). This scheme converges at the same rate as the quadrature rule considered for the numerical integration [8]. In our code, we use a trapezoidal quadrature method and hence our Nyström numerical scheme is of second order. Note that, in the case where we consider random distributions of the centers of the resonators, the Monte-Carlo Nyström numerical scheme developed in [12] can be used.
Let $\left(y_{i}\right)_{i=1}^{M}$ be a set of $M$ uniformly distributed points in $\mathcal{U}$. Discretising the Lippmann-Schwinger equation yields
$\boldsymbol{\psi}(y_{i})=\boldsymbol{v}^{\mathrm{in}}(y_{i})+\frac{|\mathcal{U}|}{M}\sum\limits_{j=1}^{M}\mathcal{V}(y_{i},y_{j})\boldsymbol{\psi}(y_{j}),$ (5.1)
@@
$\boldsymbol{\psi}^{M}:=\begin{bmatrix}\boldsymbol{\psi}(y_{1})\\ \vdots\\ \boldsymbol{\psi}(y_{M})\end{bmatrix},\quad\boldsymbol{v}^{\mathrm{in},M}:=\begin{bmatrix}\boldsymbol{v}^{\mathrm{in}}(y_{1})\\ \vdots\\ \boldsymbol{v}^{\mathrm{in}}(y_{M})\end{bmatrix},\quad\mathbb{V}:=\left(\mathcal{V}(y_{i},y_{j})\right)_{i,j=1}^{M}.$ (5.2)
@@
$\left(I_{M(2K+1)}-\frac{|\mathcal{U}|}{M}\mathbb{V}\right)\boldsymbol{\psi}^{M}=\boldsymbol{v}^{\mathrm{in},M}.$ (5.3)
@@
Figure 5.1: The effective field $\boldsymbol{\psi}^{M}$ (-) and the total field $u$ (*) plotted next to each other over $\mathcal{U}:=[0,4000]$ with $N=100$ resonators each of length $\ell=0.01$, thus $\Lambda=1$. We set $\gamma=0.05,\,\mu=0.11,\,\xi=0.2,\,K=4,\,\varepsilon_{\kappa}=0.4$. We evaluate the solution at ten time-steps in $[0,T]$, for $T=2\pi/\Omega=3141.59$. For the definition of (5.3) we choose a mesh of size $M=200$.
Figure 5.2: The $\ell^{2}$-norm of the error between the effective field $\boldsymbol{\psi}^{M}$ and the total field $u$ at time $t=0$. For the numerical solution we compare $M=N$, $M=100$ and $M=1000$. The resonators are evenly distributed inside $\mathcal{U}=[0,4000]$ and each of length $\ell=1/N$. We set $\gamma=0.05,\,\mu=0.11,\,\xi=0.2,\,K=4,\,\varepsilon_{\kappa}=0.4$.
The bold line in Figure 5.1 shows the effective field $\boldsymbol{\psi}^{M}$ as a function of $x$ evaluated at different time-steps inside the interval $[0,T]$. For comparison, we plot the evaluation of the total wave field obtained using the scheme obtained in [3], marked by stars inside the plot. As analytically proved, the effective field and the total field agree for $\ell\ll 1$ and large $N$. The numerical results in Figure 5.1 show that the effective field is still quasi-periodic with quasi-periodicity $\omega$.
Figure 5.3: The run time of the computation of $\boldsymbol{\psi}^{M}$ solving (5.3) for $M=N$, $M=100$ and $M=1000$ and $u$ solving (2.10). These results correspond to the same setting as in Figure 5.2.
Figure 5.2 shows the $\ell^{2}$-norm of the error between the effective field $\psi$ and the true solution $u$ at each resonator $z_{i}$ as a function of $N$. It becomes apparent from our numerical result that the error behaves algebraically in $N$. This result is not surprising, as the effective medium theory only holds in the limit $\ell\to 0$ and $N\to\infty$. In Figure 5.2, we specifically compare the resulting $\ell^{2}$-error for different choices of $M$. We see that the error between the effective field and the true solution decreases with increasing $N$ and $M\leq N$. Then, it is sufficient to choose $M=N$ for $N$ large enough.
The results in Figure 5.3 show how the computation run time of $u$ is significantly larger than the computation of the effective field $\boldsymbol{\psi}$ for $M=N$, $M=100$ and $M=1000$. This further underscores the relevance of the effective medium theory derived herein. Note that the run time of the computation of $\boldsymbol{\psi}^{M}$ for fixed $M(=100,1000)$ is (almost) constant in $N$.